# v-representability and Hohenberg-Kohn theorem for non-interacting Schrödinger operators with distributional potentials in the one-dimensional torus

### v1

@@ 4. Non-degeneracy of the ground-state
$|\widetilde{v}(\psi)|\lesssim\left(\lVert v\rVert_{\mathrm{H}^{-1}}+\lVert\widetilde{\varphi}\rVert_{\mathrm{H}^{1}}\right)\lVert\psi\rVert_{\mathrm{H}^{1}(\mathbb{T})},\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$,}$
@@
$-\Delta\widetilde{\varphi}(p_{1}\psi)$ $=\int_{I_{1}}\nabla\widetilde{\varphi}(x)\nabla(p_{1}\psi)(x)\mathrm{d}x=\int_{\mathbb{T}}\nabla\varphi_{v}(x)\nabla(p_{1}\psi)(x)\mathrm{d}x=v(p_{1}\psi\varphi_{v})=-v(p_{1}\psi\widetilde{\varphi}),\quad\mbox{for any$\psi\in\mathrm{H}^{1}(\mathbb{T})$,}$
$-\Delta\widetilde{\varphi}(p_{2}\psi)$ $=\int_{I_{2}}\nabla\widetilde{\varphi}(x)\nabla(p_{2}\psi)(x)\mathrm{d}x=\int_{\mathbb{T}}\nabla\varphi_{v}(x-\pi)\nabla(p_{2}\psi)(x)\mathrm{d}x=\int_{\mathbb{T}}\nabla\varphi_{v}(x)\nabla(\tau_{-\pi}p_{2}\psi)(x)$
$=-v(\varphi_{v}\tau_{-\pi}p_{2}\psi)=-v(\tau_{-\pi}p_{2}\psi\widetilde{\varphi}),\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$,}$
@@
$=-v(p_{1}\widetilde{\varphi}\psi)-v(\tau_{-\pi}p_{2}\widetilde{\varphi}\psi)+\int_{\mathbb{T}}\nabla\widetilde{\varphi}(x)\nabla\left(\frac{p_{3}}{\widetilde{\varphi}}\widetilde{\varphi}\psi\right)(x)\mathrm{d}x=-\widetilde{v}(\widetilde{\varphi}\psi),\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$.}$
Therefore $\widetilde{\varphi}$ is an eigenfunction of $h(\widetilde{v})$. Since $\widetilde{\varphi}\geq 0$ and the ground-state of $h(\widetilde{v})$ is almost everywhere strictly positive (by Lemma 4.1), the two eigenfunctions must have a non-zero overlap. In particular, they can not be orthogonal to each other. Therefore, by the non-degeneracy of the ground-state of $h(\widetilde{v})$ (see Lemma 4.1), the function $\widetilde{\varphi}$ must be the ground-state of $h(\widetilde{v})$, which yields the desired contradiction and concludes the proof. ∎
@@
Let $v\in\mathcal{V}$ and denote by $\varphi_{v}$ the ground-state of $h(v)$. Since this ground-state is non-degenerate, it follows from Lemma 3.6 that any normalized ground-state of $H_{N}(v)$ must have $\varphi_{v}$ as a natural orbital. Thus, by Theorem 2.1 and (3.9) we have
@@ 5. The non-interacting Hohenberg-Kohn theorem for distributional potentials
$-\Delta\sqrt{\rho}(\varphi)+v(\sqrt{\rho}\varphi)=0,\quad\mbox{for any $\varphi\in\mathrm{H}^{1}(\mathbb{T})$.}$
Therefore, $\sqrt{\rho}$ is an eigenfunction of $h(v)$. As $\sqrt{\rho}$ is strictly positive, it must be the ground-state of $h(v)$ by Theorem 2.1.
@@
$(\psi,T)\rightarrow h(v,T)=v\circ T\in\mathrm{H}^{-1}(\mathbb{T}),\quad\mbox{where}\quad v\circ T(\varphi)=v(T\varphi),\quad\mbox{for $\varphi\in\mathrm{H}^{1}(\mathbb{T})$,}$
@@
As $\varphi$ is strictly positive, the operator of multiplication by $\overline{\varphi}$ is an isomorphism in $\mathrm{H}^{1}(\mathbb{T})$ by Lemma 3.2. Equation (5.1) thus implies that $(v-v^{\prime})=0$, which concludes the proof. ∎
@@
As the same argument holds when we exchange $v$ and $v^{\prime}$, we conclude that $(v^{\prime}-v)(\rho)=0$. In particular, the above is an equality and $\Psi$ is also a ground-state of $H_{N}(v^{\prime})$.
Since $\Psi$ is a simultaneous ground-state of $H_{N}(v^{\prime})$ and $H_{N}(v)$, it follows from Lemma 3.6 that the ground-state of both single-particle operators $h(v)$ and $h(v^{\prime})$ is a natural orbital of $\Psi$. As all natural orbitals are mutually orthogonal, and both ground-states are strictly positive, the ground-state of $h(v)$ and $h(v^{\prime})$ must be the same. Hence, the result follows from Theorem 2.7. ∎
@@
Let $\varphi_{k}$ be a real-valued excited state of $h(v)$ for some $v\in\mathrm{H}^{-1}(\mathbb{T})$. Since the ground-state $\varphi_{1}$ is non-degenerate and strictly positive by Theorem 2.1, there must exist some $x_{0}\in T$ such that $\varphi_{k}(x_{0})=0$, as otherwise $\varphi_{k}$ would not change sign and the overlap $\langle\varphi_{k},\varphi_{1}\rangle$ would be non-zero, contradicting the orthogonality with $\varphi_{1}$. We can now consider perturbations of $v$ of the form $v_{\alpha}=v+\alpha\delta_{x_{0}}$ for $\alpha\in\mathbb{R}$. Indeed, as $\varphi_{k}$ is eigenfunction of $h(v)$ with $\varphi_{k}(x_{0})=0$, it follows that
@@
Thus $\varphi_{k}$ is also an excited state of $h(v_{\alpha})$ for any $\alpha\in\mathbb{R}$. Moreover, as the spectrum of $h(v)$ is discrete, for $\alpha>0$ small enough, $\varphi_{k}$ must be the $k^{th}$ [^7] Assuming $k-1\in\mathbb{N}$ is the previous closed shell in the case of degeneracies, i.e., $\lambda_{k-1}<\lambda_{k}$. excited state of $h(v_{\alpha})$ by standard perturbation theory. ∎
@@ 6. Concluding remarks
In this paper, we obtained a complete characterization of the set of densities that are representable by non-interacting Schrödinger operators with a large class of distributional potentials in the one-dimensional torus. Moreover, we proved a Hohenberg-Kohn theorem for such operators, thereby establishing the uniqueness of the Kohn-Sham density-to-potential map and, thanks to [SPR+24, Corollary 19], the differentiability of the non-interacting convex Lieb functional. In particular, our results show that, in the non-interacting case, the class of distributions in $\mathrm{H}^{-1}(\mathbb{T})$ is not only sufficient but also necessary to represent all strictly positive densities coming from wave-functions with finite kinetic energy.
@@
    Third, we note that Theorems 2.1 and 2.5 can be extended to the case of an interval $I=[0,2\pi]$ with Dirichlet boundary conditions. However, it is not true that all densities that are strictly positive inside $I$, and come from wave-functions with finite kinetic energy, are representable by potentials in the dual space of $\mathrm{H}^{1}_{0}(I)$. This can be seen, for instance, by considering densities like $\sqrt{\rho}(x)=x^{2}$ for $x$ close to 00, in the case of a single-particle. Indeed, in this case, $\Delta\sqrt{\rho}/\sqrt{\rho}(x)=1/x^{2}$, which does not define a continuous functional in $\mathrm{H}^{1}_{0}(I)$. Therefore, it would be interesting to systematically study the properties of the single-particle Kohn-Sham map (2.5) for functions in $\mathrm{H}^{1}_{0}(I)$. This could lead to further insights into natural necessary and sufficient conditions for $v$-representability in non-compact spaces such as the line $\mathbb{R}$, or even to higher dimensional spaces.
@@ Acknowledgements
The author is grateful to Markus Penz, Sarina Sutter, and Michael Ruggenthaler for helpful discussions, useful remarks on the first draft of this paper, and inspiring presentations during the workshop on Foundations and Extensions of DFT in Oslo. The author is also grateful to Andre Laestadius and Vebjørn Bakkestuen for organizing the aforementioned workshop which led to the ideas for the current paper.

### v2

@@ 4. Non-degeneracy of the ground-state
$|\widetilde{v}(\psi)|\lesssim\left(\lVert v\rVert_{\mathrm{H}^{-1}}+1+\frac{\lVert\widetilde{\varphi}\rVert_{\mathrm{H}^{1}}^{2}}{c^{2}}\right)\lVert\psi\rVert_{\mathrm{H}^{1}(\mathbb{T})},\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$.}$
@@
$-\Delta\widetilde{\varphi}(p_{1}\psi)$ $=\int_{I_{1}}\nabla\widetilde{\varphi}(x)\nabla(p_{1}\psi)(x)\mathrm{d}x=\int_{\mathbb{T}}\nabla\varphi_{v}(x)\nabla(p_{1}\psi)(x)\mathrm{d}x=-v(p_{1}\varphi_{v}\psi)=-v(p_{1}\widetilde{\varphi}\psi),\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$,}$
$-\Delta\widetilde{\varphi}(p_{2}\psi)$ $=\int_{I_{2}}\nabla\widetilde{\varphi}(x)\nabla(p_{2}\psi)(x)\mathrm{d}x=\int_{\mathbb{T}}\nabla\varphi_{v}(x-\pi)\nabla(p_{2}\psi)(x)\mathrm{d}x=\int_{\mathbb{T}}\nabla\varphi_{v}(x)\nabla(\tau_{\pi}p_{2}\psi)(x)$
$=-v\left(\varphi_{v}\tau_{\pi}(p_{2}\psi)\right)=-v\left(\tau_{\pi}(p_{2}\widetilde{\varphi}\psi)\right),\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$,}$
@@
$=-v(p_{1}\widetilde{\varphi}\psi)-v\left(\tau_{\pi}(p_{2}\widetilde{\varphi}\psi)\right)+\int_{\mathbb{T}}\nabla\widetilde{\varphi}(x)\nabla\left(\frac{p_{3}\widetilde{\varphi}\psi}{\widetilde{\varphi}}\right)(x)\mathrm{d}x=-\widetilde{v}(\widetilde{\varphi}\psi),\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$.}$
Therefore $\widetilde{\varphi}$ is an eigenfunction of $h(\widetilde{v})$. Since both $\widetilde{\varphi}$ and the ground-state of $h(\widetilde{v})$ are almost everywhere strictly positive (by Lemma 4.1), they can not be orthogonal to each other. Therefore, by the non-degeneracy of the ground-state of $h(\widetilde{v})$ (see Lemma 4.1), the function $\widetilde{\varphi}$ must be the ground-state of $h(\widetilde{v})$, which yields the desired contradiction and concludes the proof. ∎
@@
Let $v\in\mathcal{V}$ and denote by $\varphi_{v}$ the ground-state of $h(v)$. Since this ground-state is non-degenerate, it follows from Lemma 3.8 that any normalized ground-state of $H_{N}(v)$ must have $\varphi_{v}$ as a natural orbital with occupation number $1$. Thus, by Theorem 2.1 and (3.11) we have
@@ 5. The non-interacting Hohenberg-Kohn theorem for distributional potentials
$-\Delta\sqrt{\rho}(\varphi)+v(\sqrt{\rho}\varphi)=0,\quad\mbox{for any $\varphi\in\mathrm{H}^{1}(\mathbb{T})$, where $v=v^{\rm KS}_{1}(\rho)$.}$
Therefore, $\sqrt{\rho}$ is an eigenfunction of $h(v)$. As $\sqrt{\rho}$ is strictly positive, it must be the ground-state of $h(v)$ by Theorem 2.1. Hence, $v^{\rm KS}_{1}$ is indeed the Kohn-Sham map.
@@
$(v,T)\rightarrow h(v,T)=v\circ T\in\mathrm{H}^{-1}(\mathbb{T}),\quad\mbox{where}\quad v\circ T(\varphi)=v(T\varphi),\quad\mbox{for $\varphi\in\mathrm{H}^{1}(\mathbb{T})$,}$
@@
As $\varphi$ is strictly positive, the operator of multiplication by $\overline{\varphi}$ is an isomorphism in $\mathrm{H}^{1}(\mathbb{T})$ by Lemma 3.2. Equation (5.1) thus implies that $(v-v^{\prime})=(v-v^{\prime})\circ M_{\overline{\varphi}}\circ M_{1/\overline{\varphi}}=0$, which concludes the proof. ∎
@@
As the same argument holds when we exchange $v$ and $v^{\prime}$, we conclude that $(v^{\prime}-v)(\rho)=0$. In particular, the above is an equality and $\Psi$ is also a ground-state of $H_{N}(v^{\prime})$. Moreover, we note that, if $\rho$ is only representable by a mixed ground-state, i.e., $\rho=\sum_{j}\lambda_{j}\rho_{\Psi_{j}}$ for some $\lambda_{j}\geq 0$ and $\Psi_{j}$ ground-states of $H_{N}(v)$, the same argument shows that each of the $\Psi_{j}$ is a mutual ground-state of $H_{N}(v)$ and $H_{N}(v^{\prime})$ (see, e.g., [PTC+23, Theorem 1] for more details). So without loss of generality, we denote by $\Psi$ (one of) the mutual ground-states of $H_{N}(v)$ and $H_{N}(v^{\prime})$.
Since $\Psi$ is a simultaneous ground-state of both $H_{N}(v^{\prime})$ and $H_{N}(v)$, we must have
$\langle\left(H_{N}(v)-H_{N}(v^{\prime})\right)\Psi,\Phi\rangle=(v-v^{\prime})(\rho_{\Psi\Phi})=0,\quad\mbox{for any $\Phi\in\mathcal{H}_{N}\cap\mathrm{H}^{1}(\mathbb{T}^{N})$,}$
where $\rho_{\Psi\Phi}$ is the overlapping density
$\rho_{\Psi\Phi}(x)=N\int_{\mathbb{T}^{N-1}}\overline{\Psi(x,x_{2},...,x_{N})}\Phi(x,x_{2},...,x_{N})\mathrm{d}x_{2}...\mathrm{d}x_{N}.$
As $\Psi$ is the ground-state of $H_{N}(v)$ and $h(v)$ has discrete spectrum, by Lemma 3.8 we can choose a basis of eigenfunctions $\{\varphi_{j}\}_{j\in\mathbb{N}}$ such that $\{\varphi_{j}\}_{j\leq M}$ for some $M<\infty$ are the natural orbitals (with non-zero occupation number) of $\Psi$. Thus, if we choose trial states of the form $\Phi=a(\varphi_{\ell})^{\ast}a(\varphi_{k})\Psi$ for $k\leq M$ and $\ell\geq M+1$, from the formula in Lemma 3.9 we have
$n_{k}(v-v^{\prime})(\overline{\varphi}_{k}\varphi_{\ell})=0,\quad\quad\mbox{for any $\ell\geq M+1$ and $1\leq k\leq M$}$ (5.2)
We can now use the representation in Lemma 3.5 to show that $(v-v^{\prime})$ is more regular than expected. Precisely, note that the operator $M_{\overline{\varphi_{k}}}$ of multiplication by $\overline{\varphi_{k}}$ is a bounded operator in $\mathrm{H}^{1}(\mathbb{T})$ (see Lemma 3.2). Therefore, it follows from (5.2), the fact that $n_{j}>0$ for any $j\leq M$, and the representation in Lemma 3.5 that
$(v-v^{\prime})\circ M_{\overline{\varphi_{k}}}(\psi)=\sum_{j=1}^{M}A_{kj}\langle\varphi_{j},\psi\rangle\quad\mbox{for some $\{A_{kj}\}_{k,j\leq M}\in\mathbb{C}^{M\times M}$ and $\psi\in\mathrm{H}^{1}(\mathbb{T})$. }$ (5.3)
As the ground-state $\varphi_{1}$ is strictly positive by Theorem 2.1, the operator $M_{1/\varphi_{1}}$ is bounded in $\mathrm{H}^{1}(\mathbb{T})$ by Lemma 3.2; hence,
$(v-v^{\prime})(\psi)=(v-v^{\prime})\circ M_{\overline{\varphi_{1}}}\circ M_{1/\overline{\varphi_{1}}}(\psi)=\sum_{j=1}^{M}A_{1j}\left\langle\frac{\varphi_{j}}{\varphi_{1}},\psi\right\rangle,\quad\mbox{for any $\psi\in\mathrm{H}^{1}(\mathbb{T})$.}$
In other words, the distribution $(v-v^{\prime})$ can be identified (via the Riesz map (3.3)) with the function
$(v-v^{\prime})(x)=\sum_{j=1}^{M}A_{1j}\frac{\varphi_{j}(x)}{\varphi_{1}(x)}.$ (5.4)
In particular $(v-v^{\prime})\in\mathrm{H}^{1}(\mathbb{T})$. Hence, to conclude the proof, it suffices to show that $(v-v^{\prime})$ is piecewise constant. Indeed, if $(v-v^{\prime})$ was piecewise constant but not constant, it would have a jump, which is not possible for $\mathrm{H}^{1}$ functions.
To see that $(v-v^{\prime})$ is piecewise constant, note that by (5.3) we have
$(v-v^{\prime})(x)\frac{\varphi_{k}(x)}{\varphi_{1}(x)}=\sum_{j=1}^{N}A_{kj}\frac{\varphi_{j}(x)}{\varphi_{1}(x)},\quad\mbox{for any $1\leq k\leq M$ and (almost) every $x\in\mathbb{T}$.}$
In other words, the above equation shows that the vector
$\vec{\Phi}(x)\coloneqq\left(1,\frac{\varphi_{2}(x)}{\varphi_{1}(x)},...,\frac{\varphi_{M}(x)}{\varphi_{1}(x)}\right)^{T}\neq 0$
is an eigenvector of the matrix $A=\{A_{kj}\}\in\mathbb{C}^{M\times M}$ with corresponding eigenvalue $(v-v^{\prime})(x)$ for (almost) every $x\in\mathbb{T}$. As $A\in\mathbb{C}^{M\times M}$ can have at most $M$ distinct eigenvalues, we conclude that $(v-v^{\prime})(x)$ is piecewise constant, which completes the proof.∎
@@
Let $\varphi_{k}$ be a real-valued excited state of $h(v)$ for some $v\in\mathrm{H}^{-1}(\mathbb{T})$. Since the ground-state $\varphi_{1}$ is non-degenerate and strictly positive by Theorem 2.1, there must exist some $x_{0}\in\mathbb{T}$ such that $\varphi_{k}(x_{0})=0$, as otherwise $\varphi_{k}$ would not change sign and the overlap $\langle\varphi_{k},\varphi_{1}\rangle$ would be non-zero, contradicting the orthogonality with $\varphi_{1}$. We can now consider perturbations of $v$ of the form $v_{\alpha}=v+\alpha\delta_{x_{0}}$ for $\alpha\in\mathbb{R}$. Indeed, as $\varphi_{k}$ is an eigenfunction of $h(v)$ with $\varphi_{k}(x_{0})=0$, it follows that
@@
Thus $\varphi_{k}$ is also an excited state of $h(v_{\alpha})$ for any $\alpha\in\mathbb{R}$. Moreover, as the spectrum of $h(v)$ is discrete, for $\alpha>0$ small enough, $\varphi_{k}$ must be the $k^{th}$ excited state [^8] Assuming $k-1\in\mathbb{N}$ is the previous closed shell in the case of degeneracies, i.e., $\lambda_{k-1}<\lambda_{k}$. of $h(v_{\alpha})$ by standard perturbation theory. ∎
@@ 6. Concluding remarks
In this paper, we obtained a complete characterization of the set of densities that are representable by non-interacting Schrödinger operators with a certain class of distributional potentials in the one-dimensional torus. Moreover, we proved a Hohenberg-Kohn theorem for such operators, thereby establishing the uniqueness of the Kohn-Sham density-to-potential map and, thanks to [SPR+24, Corollary 19], the differentiability of the non-interacting convex Lieb functional. In particular, our results show that, in the non-interacting case, the class of distributions in $\mathrm{H}^{-1}(\mathbb{T})$ is not only sufficient but also necessary to represent all strictly positive densities coming from wave-functions with finite kinetic energy.
@@
    Third, we note that Theorems 2.1 and 2.5 can be extended to the case of an interval $I=[0,2\pi]$ with Dirichlet boundary conditions. However, it is not true that all densities that are strictly positive inside $I$, and come from wave-functions with finite kinetic energy, are representable by potentials in the dual space of $\mathrm{H}^{1}_{0}(I)$. This can be seen, for instance, by considering densities like $\sqrt{\rho}(x)=x^{2}$ for $x$ close to 00, in the case of a single-particle. Indeed, in this case, $\Delta\sqrt{\rho}/\sqrt{\rho}(x)=2/x^{2}$, which does not define a continuous functional in $\mathrm{H}^{1}_{0}(I)$. Therefore, it would be interesting to systematically study the properties of the single-particle Kohn-Sham map (2.5) for functions in $\mathrm{H}^{1}_{0}(I)$. This could lead to further insights into natural necessary and sufficient conditions for $v$-representability in non-compact spaces such as the line $\mathbb{R}$, or even to higher dimensional spaces.
@@ Acknowledgements
The author is grateful to Markus Penz, Sarina Sutter, and Michael Ruggenthaler for helpful discussions, useful remarks on the first draft of this paper, and inspiring presentations during the workshop on Foundations and Extensions of DFT in Oslo. The author is also grateful to Andre Laestadius and Vebjørn Bakkestuen for organizing the aforementioned workshop which led to the ideas for the current paper. Special thanks also to Asbjørn Bækgaard Lauritsen for calling attention to Remark 3.7 and spotting a mistake in the previous proof of Theorem 2.5.